# Digital Durable Goods Monopoly

### v8

@@ 5.1 Positive Marginal Cost
Our results indicate that within the Coasian bargaining framework, optimal market-clearing profit and market efficiency (or the Folk Theorem) constitute distinct economic forces. The first, identified by Nava and Schiraldi (2019), derives from intratemporal price discrimination. The second, which is the focus of this analysis, stems from intertemporal price discrimination.
@@
Figure 1: Conceptual Relationship to the Standard Coasian Model
@@
Lastly, our work relates to the literature on reputational bargaining, initiated by Abreu and Gul (2000). While that literature rationalizes delay and inefficiency through the presence of behavioral types, we do not model commitment as an exogenous characteristic. Instead, we identify reputational effects that arise endogenously from the structural properties of digital goods.
@@
Our results show how specific technological features alter the strategic structure of dynamic monopoly. When allocative efficiency is set-valued rather than single-valued such as in digital goods markets, Coasian forces do not generally select a unique outcome. Given the indeterminacy, realized outcomes may depend on norms, reputation, or seller-specific characteristics rather than solely on fundamentals such as demand, buyer valuations, or production costs.
This implication is consistent with observed patterns in the B2B digital goods sector, as illustrated by Broadcom’s acquisition of VMware. VMware’s core product, vSphere, is a server virtualization platform, a software layer that partitions a physical server into multiple virtual machines. As documented, Broadcom terminated existing licenses and required customers to renegotiate, resulting in revised terms, with renewal prices increasing by 800% and, in some cases, exceeding 1,500%  [^13] See https://www.ciodive.com/news/broadcom-vmware-lock-in-cost-disruption/718281/.. Although the seller changed from VMware to Broadcom, the product itself did not change.
This raises the relevance of regulatory oversight. Although optimal policy design lies beyond the scope of this paper, the analysis suggests that market forces alone may not generate consistent or efficient outcomes.
@@ A.1 Proofs of the Results in Section 3
The right-hand side equals the expected payoff for type $\theta$ when mimicking type $\theta^{\prime}$ in the dynamic game. Since $\alpha$ is a best response for the buyer, the induced static mechanism satisfies incentive compatibility and individual rationality. By construction, the seller’s expected payoff in the static mechanism coincides with her payoff under $(\sigma,\alpha)$ in the dynamic game. Therefore, every dynamic payoff can be replicated in the static setting, which implies that the optimal static payoff is an upper bound on the seller’s achievable payoffs.
@@ A.2 Proofs of the Results in Section 4
    By construction, type $\theta(\delta,x,p)$ is indifferent between purchasing immediately at price $p$ and waiting to purchase the allocation $\max\{x,x^{e}(\underline{\theta})\}$ in the subsequent period at price $u_{\min}$. If a solution exists, it is unique by the single-crossing property of the utility function. Although the net utility $u(x,\theta)-p$ need not satisfy single crossing in $\theta$ for a fixed offer $(x,p)$ in general, our construction of the continuation allocation $\max\{x,x^{e}(\underline{\theta})\}$ ensures that the indifference condition exhibits single crossing. In particular, the difference between the current-period payoff and the discounted continuation payoff crosses zero at most once in $\theta$, which guarantees uniqueness of the cutoff type.
@@
However, by Lemma 7, immediate market clearing in period t+n1t+n-1 at price $p=u_{\min}$ yields a strictly higher payoff than delaying market clearing to period $t+n$. Thus, the deviation is strictly dominated by immediate market clearing in period t+n1t+n-1. By backward induction, the same argument applies to all preceding deviation periods. We therefore conclude that no profitable deviation exists, and $(\sigma,\alpha)$ constitutes an equilibrium for all $0\leq\delta<1$.
@@
which is strictly positive and bounded away from zero as $\delta\to 1$. Conversely, the gain from accelerating sales is of order $(1-\delta)K$, where $K$ is a bounded constant representing the future surplus to be collected. This term converges to zero as $\delta\to 1$. Therefore, for $\delta$ sufficiently close to $1$, the potential gain from deviation is strictly dominated by the immediate revenue loss. Under this condition, the seller has no incentive to deviate, which completes the argument.
@@ A.3 Proof of the Results in Section 5
To complete the proof, note that the prior distribution contains a point mass at $\underline{\theta}$. Since the remaining mass $P(h_{t})$ cannot decay geometrically below the mass of this atom while the atom remains active, the process must eventually reach a state in which the remaining mass consists predominantly of type $\underline{\theta}$. At that point, the seller clears the market immediately.
@@
Normalizing by the factor $(1-\delta)$, the optimality condition implies that a positive mass of buyers must clear in each period. This requirement yields a bound $T$ that is independent of $\delta$. Consequently, as $\delta\to 1$, the discount factor over the entire horizon, $\delta^{T}$, converges to $1$. The resulting outcome combines efficient allocations with vanishing delay.
@@ B.1 Payoff Upper Bound
In this section, we show that $\pi^{e}(F)$ is an upper bound on the seller’s payoff within a broad class of equilibria. As shown above, when $x^{e}(\underline{\theta})\leq\underline{x}$, we have $\pi^{e}(F)=\pi(F)$. This case is comparable to Ausubel and Deneckere (1989b), where the lowest type has zero valuation. In both settings, the lowest type is indifferent across all feasible allocation outcomes, which in the setting of Ausubel and Deneckere (1989b) corresponds to the discounted probability of trade.
We therefore focus on the case in which $x^{e}(\underline{\theta})>\underline{x}$. We begin by introducing the following definitions.
@@ B.2 Relaxing Assumption 3
Using the equilibrium outcome characterized in Proposition 3 as the off-path punishment, we can sustain the menu described in Lemma 4. The proof is analogous to that of Theorem 2 and is therefore omitted. Since the action space is discretized, the continuum may not be attainable for arbitrary $\Delta>0$.
B.3 Menu Offers
Corollary 7.
By contrast, the baseline model rules out intratemporal discrimination. In this case, the maximal market-clearing profit is $u_{\min}$, which is strictly below the static commitment benchmark. This difference explains why the equilibrium outcomes and underlying economic forces in the present model differ from those in the existing literature.
B.4 Indeterminacy of Delay
@@
B.5 Payoff Upper Bound without Free Disposability
@@
Appendix C Institutional Evidence
C.1 Importance of B2B Digital Goods Market
The market for Business-to-Business (B2B) digital goods—including enterprise software, financial data, and AI models—is a central component of the modern economy. Market leaders in this sector exhibit high profitability (Guellec and Paunov, 2017). Direct evidence comes from the disclosures of publicly traded corporations. Specifically, annual Form 10-K filings provide an audited decomposition of revenue streams.
An examination of filings from industry leaders shows a consistent pattern: the enterprise sector accounts for a large share of profits. Microsoft Corporation’s enterprise-focused segments constitute the primary drivers of its profitability. The Intelligent Cloud segment generated \$87.9 billion in revenue in fiscal year 2023. The Productivity and Business Processes segment derives most of its revenue from commercial clients, with Office Commercial products and cloud services contributing \$48.7 billion [^17] See Microsoft Corporation (2023).. Salesforce reported \$34.9 billion in revenue for fiscal year 2024, generated from B2B subscriptions to its cloud-based software [^18] See Salesforce, Inc. (2024).. Adobe Inc.’s Digital Media segment generated \$14.2 billion in 2023, a substantial portion of which came from enterprise subscriptions. The Digital Experience segment is an exclusively B2B division providing marketing, analytics, and e-commerce services, generating \$4.9 billion in revenue [^19] See Adobe Inc. (2023)..
This underscores the importance of understanding the market dynamics in this industry.
C.2 Single-offer Protocol
The model assumes that the seller makes a single take-it-or-leave-it offer in each period. Although the confidential nature of B2B negotiations often precludes direct observation of the bargaining process, we discuss the Amazon Web Services “Private Offer” feature as an institutional commitment device.
The institutional design of major B2B marketplaces, including the AWS Marketplace, is consistent the single-offer protocol. On these platforms, Independent Software Vendors (ISVs) such as Snowflake and data providers such as Reuters transact with enterprise buyers. To facilitate these transactions, the marketplace employs a “Private Offer” mechanism (Amazon Web Services (2025)).
This mechanism serves as an institutional commitment device. The seller issues a formal, binding proposal to a specific buyer that is valid for a fixed duration. If the buyer rejects the offer, the seller may issue a subsequent proposal. Crucially, each offer consists of a single, indivisible set of terms, including the software tier, user count, and price. The buyer therefore faces a binary acceptance decision over a bundled proposal. This design is consistent with the model’s structural assumption: regardless of informal pre-negotiation, the platform’s architecture requires the seller to commit to one binding offer at a time. The fact that the industry-standard instrument for executing customized transactions takes the form of a single offer is consistent with for the framework.

### v10

@@ 5.3 Menu Offers
The Coasian durable-goods literature studies monopoly pricing under limited commitment and typically focuses on environments in which allocative efficiency is single-valued, such as timing or probabilistic trade. In these settings, efficiency tightly restricts continuation play and leads to sharp selection results in the patient limit. By contrast, information goods, in which delivered scope or precision can exceed efficient utilization once usage is chosen ex post, have not been studied in this literature. This paper fills that gap by analyzing Coasian dynamics in environments with set-valued efficiency.
@@
Our results indicate that within the Coasian bargaining framework, optimal market-clearing profit and allocative efficiency (at least for the lowest type) constitute distinct economic forces. The first, identified by Nava and Schiraldi (2019), derives from intratemporal price discrimination. The second, which is the focus of this analysis, stems from intertemporal price discrimination.
@@
Figure 1: Relationship to the Standard Coasian Model
@@
Lastly, our work relates to the literature on reputational bargaining, initiated by Abreu and Gul (2000). While that literature rationalizes delay and inefficiency through the presence of behavioral types, we do not model commitment as an exogenous characteristic. Instead, we identify reputational effects that arise endogenously from the set-valued allocative efficiency in our settings.
@@
This paper studies dynamic monopoly under limited commitment in environments in which allocative efficiency is set-valued rather than single-valued. We show that Coasian discipline continues to constrain outcomes through the lowest-type buyer’s allocation but does not generally select a unique equilibrium path. As a result, the set of equilibrium seller payoffs can exhibit substantial indeterminacy in the patient limit.
Information goods provide a benchmark for this analysis. Interpreted through this lens, the results shift attention from point predictions to the structure of the equilibrium payoff set under limited commitment.
@@ A.1 Proofs of the Results in Section 3
The right-hand side equals the expected payoff for type $\theta$ when mimicking type $\theta^{\prime}$ in the dynamic game. Since $\alpha$ is a best response for the buyer, the induced static mechanism satisfies incentive compatibility and individual rationality. By construction, the seller’s expected payoff in the static mechanism coincides with her payoff under $(\sigma,\alpha)$ in the dynamic game. Therefore, every dynamic payoff can be replicated in the static setting, which implies that the optimal static payoff is an upper bound on the seller’s achievable payoffs. ∎
@@ A.2 Proofs of the Results in Section 4
    By construction, type $\theta(\delta,x,p)$ is indifferent between purchasing immediately at price $p$ and waiting to purchase the allocation $\max\{x,x^{e}(\underline{\theta})\}$ in the subsequent period at price $u_{\min}$. If a solution exists, it is unique by the single-crossing property of the utility function.
    Although the net utility $u(x,\theta)-p$ need not satisfy single crossing in $\theta$ for a fixed offer $(x,p)$ in general, our construction of the continuation allocation $\max\{x,x^{e}(\underline{\theta})\}$ ensures that the indifference condition exhibits single crossing. In particular, the difference between the current-period payoff and the discounted continuation payoff crosses zero at most once in $\theta$, which guarantees uniqueness of the cutoff type.
@@
However, by Lemma 7, immediate market clearing in period t+n1t+n-1 at price $p=u_{\min}$ yields a strictly higher payoff than delaying market clearing to period $t+n$. Thus, the deviation is strictly dominated by immediate market clearing in period t+n1t+n-1. By backward induction, the same argument applies to all preceding deviation periods. We therefore conclude that no profitable deviation exists, and $(\sigma,\alpha)$ constitutes an equilibrium for all $0\leq\delta<1$. ∎
@@
which is strictly positive and bounded away from zero as $\delta\to 1$.
Conversely, the gain from accelerating sales is of order $(1-\delta)K$, where $K$ is a bounded constant representing the future surplus to be collected. This term converges to zero as $\delta\to 1$. Therefore, for $\delta$ sufficiently close to $1$, the potential gain from deviation is strictly dominated by the immediate revenue loss. Under this condition, the seller has no incentive to deviate, which completes the argument. ∎
@@ A.3 Proof of the Results in Section 5
To complete the proof, note that the prior distribution contains a point mass at $\underline{\theta}$. Since the remaining mass $P(h_{t})$ cannot decay geometrically below the mass of this atom while the atom remains active, the process must eventually reach a state in which the remaining mass consists predominantly of type $\underline{\theta}$. At that point, the seller clears the market immediately. ∎
@@
Normalizing by the factor $(1-\delta)$, the optimality condition implies that a positive mass of buyers must clear in each period. This requirement yields a bound $T$ that is independent of $\delta$. Consequently, as $\delta\to 1$, the discount factor over the entire horizon, $\delta^{T}$, converges to $1$. The resulting outcome combines efficient allocations with vanishing delay. ∎
@@ B.1 Payoff Upper Bound
In this section, we show that $\pi^{e}(F)$ is an upper bound on the seller’s payoff within a broad class of equilibria. As shown above, when $x^{e}(\underline{\theta})\leq\underline{x}$, we have $\pi^{e}(F)=\pi(F)$. We therefore focus on the case in which $x^{e}(\underline{\theta})>\underline{x}$. We begin by introducing the following definitions.
@@ B.2 Relaxing Assumption 3
Using the equilibrium outcome characterized in Proposition 3 as the off-path punishment, we can sustain the menu described in Lemma 4. The proof is analogous to that of Theorem 2 and is therefore omitted. Since the action space is discretized, the continuum may not be attainable for arbitrary $\Delta>0$ [^11] Nonetheless, the payoff set can converge to the continuum as $\Delta\rightarrow 0$..
B.3 Indeterminacy of Delay
@@
B.4 Payoff Upper Bound without Free Disposability
@@
Appendix C Single-offer Protocol
The model assumes that the seller makes a single take-it-or-leave-it offer in each period. Although the confidential nature often precludes direct observation of the bargaining process, we discuss the Amazon Web Services “Private Offer” feature as an institutional example.
On AWS Marketplace, data providers such as Reuters trade with enterprise buyers. The marketplace employs a “Private Offer” protocol (Amazon Web Services (2025)) which serves as a commitment device. The seller issues a formal, binding proposal to a specific buyer that is valid for a fixed duration. If the buyer rejects the offer, the seller may issue a subsequent proposal. Crucially, each offer consists of a single, indivisible set of terms. The buyer therefore faces a binary acceptance decision over a bundled proposal. This design is consistent with the model’s structural assumption: regardless of informal pre-negotiation, the platform’s architecture requires the seller to commit to one binding offer at a time.